1)Find the present value of $13,500 in 7 months at 7.9% interest

answer:$

2. Your friend needs a loan of $4,700 for 6 months. You tell him for the simple interest loan service he needs to pay you back $4,770.50. What is the simple interest rate on your friendly loan?

Interest Rate= %


3.A payday loan company charges a $85 fee for a $400 payday loan that will be repaid in 13 days. Treating the fee as interest paid, what is the equivalent annual interest rate?

answer: % interest
Answer #1

1. PV = FV/(1 + r)^n

Monthly interest rate = 7.9%/12 = 0.006583333333

PV = 13,500/(1 + 0.006583333333)^7

PV = $12,893.939317269

2. Simple interest for 6 months = (Repayment amount - loan amount)/Loan mount

Simple interest rate for 6 months = (4,770.50 - 4,700)/4,700

Simple interest rate for 6 months = 0.015

Simple interest rate for 6 months = 1.5%

Or 1.5% * 2 = 3% per year

Please do not downvote for not answering the remaining questions. As per HOMEWORKLIB RULES, when there are multiple questions, we are encouraged to provide a solution to at least the first question.

Can you please upvote? Thank You :-)